The function is defined by f(x) = ax + b for x ER, where a and b are constants. It is given that f(2)= 1 and f(5) = 7. ii) Solve them and find the values of a and b. 1) Solve the equation f(f(x)) = 0. 1) Set up a pair of simultaneous equations using the information given.


1
Expert's answer
2021-11-23T08:38:01-0500


ii)

"1=a(2)+b"

"7=a(5)+b"

"a=slope=\dfrac{7-1}{5-2}=2"

"b=1-2(2)=-3"

"a=2, b=-3, y=2x-3"

1)


"f(f(x))=0"

"a(ax+b)+b=0"

"a^2x+ab+b=0"

"x=-\dfrac{ab+b}{a^2}"

"x=-\dfrac{2(-3)-3}{(2)^2}"

"x=\dfrac{9}{4}"